### v2

# Minerva: A File-Based Ransomware Detector

Abstract.
Ransomware attacks have caused billions of dollars in damages in recent years, and are expected to cause billions more in the future. Consequently, significant effort has been devoted to ransomware detection and mitigation. Behavioral-based ransomware detection approaches have garnered considerable attention recently. These behavioral detectors typically rely on process-based behavioral profiles to identify malicious behaviors. However, with an increasing body of literature highlighting the vulnerability of such approaches to evasion attacks, a comprehensive solution to the ransomware problem remains elusive.
This paper presents Minerva, a novel robust approach to ransomware detection. Minerva is engineered to be robust by design against evasion attacks, with architectural and feature selection choices informed by their resilience to adversarial manipulation. We conduct a comprehensive analysis of Minerva across a diverse spectrum of ransomware types, encompassing unseen ransomware as well as variants designed specifically to evade Minerva. Our evaluation showcases the ability of Minerva to accurately identify ransomware, generalize to unseen threats, and withstand evasion attacks. Furthermore, Minerva achieves remarkably low detection times, enabling the adoption of data loss prevention techniques with near-zero overhead.
ransomware detection, behavioral classification, machine learning
ccs: Security and privacy Malware and its mitigationccs: Computing methodologies Machine learning
1. Introduction
Despite intense research (Moussaileb et al., 2021), ransomware attacks continue to pose significant challenges. In ransomware attacks, a malicious actor infiltrates the systems of a victim organization (such as an enterprise (Derek Kortepeter, 2018) or a city government (Blinder and Perlroth, 2018)) and installs specialized malware—ransomware—that encrypts sensitive organization data. Subsequently, a ransom demand is made in exchange for decryption keys. Ransomware attacks can result in substantial financial losses for victims and disrupt critical systems, as evidenced by the recent Colonial Pipeline incident (Morgan, 2019; Ventures, 2023).
The detection of ransomware on servers and user machines is a crucial last line of defense. Several detection algorithms can be employed to identify ransomware activity, terminate unwanted processes, and restore encrypted files to their original content (Sheen and Yadav, 2018; Yu et al., 2024; Baek et al., 2018). Many existing approaches (Continella et al., 2016; Mehnaz et al., 2018; Kharraz and Kirda, 2017; Elkhail et al., 2023) leverage process-level features to differentiate ransomware processes performing file encryption from other benign processes executing on the machine. The most effective process-based detectors typically utilize behavioral detection: they construct behavioral profiles of running processes based on their system activity, and compare these profiles against a learned behavioral model of benign and ransomware processes. This approach is reasonable: ransomware operations involve intense disk I/O and encryption, typically resulting in distinct behavioral profiles compared to benign applications. Generally, these behavioral-based detectors rely on Machine Learning (ML) techniques to construct and compare behavioral profiles. ML techniques have proven to be valuable tools across various cybersecurity domains (Mehnaz and Bertino, 2021; Piskozub et al., 2021; Yin et al., 2023), yielding substantial improvements over prior art (Hang et al., 2023; Hitaj et al., 2021; Pagnotta et al., 2022). However, recent research (De Gaspari et al., 2022a; Zhou et al., 2023) has revealed a series of evasive techniques that can readily bypass both academic and commercial process-based behavioral detectors. These evasive ransomware attacks entail distributing the ransomware workload across several cooperating processes, each emulating, feature-wise, the behavior of a benign process. Effectively, the behavioral profile of each ransomware process is indistinguishable from that of benign processes, but their coordinated action still results in the expected ransomware behavior. This evasive ransomware approach exploits inherent weaknesses of process-level features and, as such, cannot be identified at the process level. These attacks fall under a broader category of adversarial attacks against ML systems (Rosenberg et al., 2021; Bundt et al., 2023; Hitaj et al., 2022), which highlight the risks associated with such approaches.
This paper introduces Minerva, a novel defense against ransomware. Unlike prior methods that primarily rely on process-level behavioral analysis, Minerva classifies system activity by constructing file-level behavioral profiles for each file, based on the operations they undergo within a defined time window. Ransomware processes are identified based on their interaction with a file during a time window in which malicious activity is detected. We also engineer Minerva to be robust by design against complex evasion attacks by basing its architecture on a contrastive design, where altering one aspect of the behavioral profile triggers detectable changes to other aspects. Minerva’s architecture is based on two key insights. (1) The first insight is that while evasive ransomware attacks may mimic benign applications at the level of individual ransomware processes, the ultimate objective of ransomware is to encrypt user files. Therefore, regardless of how tasks are distributed among these individual processes, the ransomware’s activity becomes evident from a file-based perspective. (2) The second insight recognizes that different aspects of file-based behavioral profiles are interconnected and, due to constraints in the ransomware goals, changes to one aspect directly influence others. For instance, ransomware can manipulate the file-based behavioral profile of a file by altering the average entropy of file operations. However, to fully encrypt the file, it must eventually overwrite it with high-entropy data. To satisfy this constraint, ransomware must increase the number of operations on the file, introducing low-entropy operations to decrease average entropy, thereby modifying another aspect of the behavioral profile. Figure 1 provides a high-level overview of Minerva’s detection module.
Figure 1. Overview of Minerva’s ransomware detection module. For each individual file, Minerva maintains a record of operations performed by various processes P across different temporal resolutions (windows). At the conclusion of each window, file operations are aggregated and submitted to the relevant ML classifier trained for that temporal resolution.
Summarizing, this paper makes the following contributions:
  1. (1)
    We propose Minerva, a robust ransomware detector based on the novel concepts of file-based behavioral profiles and contrastive design. Minerva overcomes critical limitations of existing behavioral detection techniques.
  2. (2)
    We perform a comprehensive analysis of the features generated by ransomware activity, identifying a robust set of features intrinsically related to core ransomware operations.
  3. (3)
    We assess Minerva’s performance against traditional, evasive multiprocess, and unseen ransomware. Our findings demonstrate that Minerva effectively detects ransomware activity on average within $0.52$ seconds of the onset of malicious activity.
  4. (4)
    We analyze the resilience of Minerva to adaptive, unseen ransomware: ransomware specifically engineered to evade Minerva’s detection. We introduce three distinct families of adaptive ransomware and validate Minerva’s robustness against each, emphasizing its capacity for generalization and underscoring the significance of our contrastive design.
  5. (5)
    We study the explainability of Minerva and provide key insights on its effectiveness.
2. Related Work
Ransomware detection approaches rely on the insight that ransomware processes behave very differently from benign processes. Such approaches commonly consist of three main components. A monitoring component monitors the dynamic behavior of running processes, which is used to compute a set of features based on program execution. Feature computation typically considers a sequence of recent process system calls whose characteristics are aggregated over a window of operations. Typical features used for this purpose are frequency and characteristics of file read/write operations, changes to file types (such as file type and file content entropy), and others. All these features are computed per process and then fed to a machine learning classifier, which uses them to label processes as benign or ransomware. Several techniques proposed in recent years follow the approach above: UNVEIL (Kharaz et al., 2016) and Redemption (Kharraz and Kirda, 2017) generate a suspicious activity score based on characteristics of file write operations, entropy changes, and others. CryptoDrop (Scaife et al., 2016) generates a reputation score computed similarly. ShieldFS (Continella et al., 2016) uses relative frequencies of various disk operations, such as read, write, folder listings, and others, as features fed to a random forest process classifier. A similar approach is also used by RWGuard (Mehnaz et al., 2018). While detection techniques used in commercial detectors are not disclosed, De Gaspari et al. (De Gaspari et al., 2020) show that Malwarebytes’s (Malwarebytes, 2024) ransomware detector is vulnerable to evasion attacks designed against process-based detectors and thus likely to leverage similar techniques.
Recently, the same authors proposed a novel family of evasion attacks against process-based ransomware detectors (De Gaspari et al., 2022a). These attacks fall under a broader category of adversarial attacks against ML systems (Rosenberg et al., 2021; De Gaspari et al., 2024; Hitaj et al., 2024), highlighting the risks associated with such approaches. The authors suggested three attacks: process splitting, functional splitting, and mimicry attack. All these attacks aim to keep behavioral features’ values below the detection threshold through various mechanisms, but the core principle is to split the ransomware operations over multiple independent processes. The process splitting attack evenly distributes the malicious activity across several processes. This approach tends to be impractical in practice, requiring the creation of many processes. The functional splitting attack spawns fewer processes, each performing a subset of ransomware operations. While this attack is practical, it can be mitigated with adversarial training (De Gaspari et al., 2022a). Finally, the mimicry attack works by splitting ransomware into multiple processes, each mimicking the behavioral profile of benign applications from the point of view of disk operations. Mimicry attacks are fundamentally hard to detect as they are designed to resemble, feature-wise, the behavior of a benign process. Ģenc et al. (Genç et al., 2018) provides a high-level discussion of similar evasive approaches. MalWASH (Ispoglou and Payer, 2016), D-TIME (Pavithran et al., 2019), and ROPE (D’Elia et al., 2021) are generic malware obfuscators that also leverage the distribution of malware workload across processes, similarly to (De Gaspari et al., 2022a). Rather than distributing malware workload, other approaches attempt to perturb the sequence of system calls generated by ransomware (Rosenberg et al., 2018; Hu and Tan, 2018).
Other approaches focus on using high entropy as a proxy to detect ransomware encryption activity. These approaches use various techniques, including $\chi^{2}$-test (Palisse et al., 2017) and the Kullback-Liebler divergence (Mbol et al., 2016). However, when used in isolation, entropy is a weak distinguisher of ransomware activity and can be easily evaded (De Gaspari et al., 2022b).
Another line of work utilizes decoy files (Genç et al., 2019; Moore, 2016; Moussaileb et al., 2018). A decoy file is a dummy file, typically hidden from the user and created for ransomware detection. As no application is expected to use a decoy file, accesses to it are a strong signal of ransomware activity. Recently, Ganfure et al. (Ganfure et al., 2023) presented RTrap, an advanced version of traditional decoy files. Rtrap is a machine learning-based approach to decoy files to provide early detection against cryptographic ransomware. In particular, RTrap selects a set of legitimate user files from the system, duplicates and deploys them throughout the system. The user files to be used as decoys are selected in such a way that they are representative of the user content on the machine. The insight behind the approach is that modern ransomware does not simply encrypt all files indiscriminately, but rather prioritizes certain types of file contents over others in order to be more effective. Panzade et al (Panzade et al., 2024) showed that RTrap can be evaded by not trying to encrypt the user files in a particular directory at the very beginning but rather, stealthily, find out which files can potentially be a decoy and then build an encryption plan that skips them.
In contrast to prior work on ransomware detection, Minerva was designed with the goal of being resilient to evasion attacks. This goal led to a detailed investigation of ransomware activity, from which we decided to monitor and model the behavior of the system activity from a file point-of-view. On top of that, the features selected for monitoring were chosen following a contrastive design paradigm, meaning that alternating one or more features would inevitably alter the rest of the remaining features, thus leaving little to no room for evasive ransomware to try to evade detection. The following section thoroughly introduces our approach and evaluates its ransomware detection capabilities. To the best of our knowledge, we are the first to fully evaluate a ransomware detection approach from an adaptive adversary’s point-of-view, a feat not done in prior ransomware detection approaches.
3. Threat Model
We position ourselves in a threat model similar to those considered by prior work in the domain (Continella et al., 2016; Mehnaz et al., 2018; Kharraz and Kirda, 2017; Elkhail et al., 2023): the adversary is a ransomware that has infected a target machine with the goal of encrypting users files and demanding a ransom to recover the data. We consider different types of ransomware, categorized into three groups according to their capabilities and behaviors:
  • Traditional Ransomware, corresponding to the typical ransomware families found in the wild.
  • Evasive Multiprocess Ransomware, corresponding to complex ransomware that mimics the behavior of benign processes to avoid detection (De Gaspari et al., 2020; Zhou et al., 2023).
  • Adaptive Ransomware, corresponding to a sophisticated ransomware engineered specifically to evade Minerva’s file-based detection through manipulation of features utilized by Minerva for detection.
Trusted Components. In our threat model we assume Minerva’s components to be trusted, meaning that the ransomware processes cannot tamper with their functionality. In particular, it is assumed that the Disk Activity Monitor module, and the File-Based Behavioral Detector module, can work correctly and independently of any ransomware infecting the system. These assumptions are in line with prior work in the field (Continella et al., 2016; Mehnaz et al., 2018; Kharraz and Kirda, 2017; Elkhail et al., 2023).
4. Minerva: A File-Based Ransomware Detector
This section presents Minerva, a novel ransomware detector built upon the concepts of file-based behavioral profiling and contrastive design. Minerva effectively guards against traditional ransomware, evasive multiprocess ransomware, and adaptive ransomware engineered specifically to evade Minerva’s detection. Minerva’s architecture is based on two key insights. (1) The first insight is that, although individual ransomware processes may modulate their behavior to imitate benign applications (De Gaspari et al., 2022a; Zhou et al., 2023), they must ultimately encrypt large portions of users files to achieve their objective. Consequently, it is possible to monitor the behavioral profiles of the files themselves and detect any deviation from expected behaviors. From a file-based standpoint, the ransomware activity becomes apparent regardless of the number of processes performing encryption, or how tasks are distributed among individual ransomware processes. Minerva leverages this insight by creating file-based behavioral profiles for each opened user file in the system, based on the I/O operations they receive in a given time window. Ransomware detection is performed per file and per time window based on the constructed profiles. (2) The second insight is that different aspects of file-based behavioral profiles are interconnected. Therefore, any attempt to alter some aspects of the behavioral profile to evade detection inevitably leads to detectable changes in others. Minerva capitalizes on this insight through its contrastive design: Minerva’s features and architecture are selected so that any manipulation by ransomware to evade detection inevitably causes changes to other features that still expose the malicious activity, or that are captured by the tiered architecture of Minerva (see Section 4.1). We provide an in-depth discussion on Minerva’s contrastive design in Section 4.2.
4.1. Minerva Architecture
From a high-level view, Minerva is comprised of two separate modules: (1) Disk Activity Monitor, and (2) File-Based Behavioral Detector. In what follows, we explain how these two modules work.
(a) R/W ent. ratio.
(b) Num. of operations.
(c) Num. of processes.
(d) R/W ratio.
(e) R/W ent. ratio CDF.
(f) Num. of ops. CDF.
(g) Num of Proc. CDF.
(h) R/W ratio CDF.
Figure 2. Comparison of distribution and Cumulative Distribution Function (CDF) of different Minerva features for benign and ransomware processes. Computed on 1 second window.
4.1.1. Disk Activity Monitor.
The Disk Activity Monitor module is a shim that provides system-wide, real-time information on all file system operations performed. We implemented the module as a driver that uses the I/O Request Packet data structure provided by Windows systems to transparently log filesystem operations system-wide. Each filesystem-related system call performed by a process can result in many lower-level operations. The Disk Activity Monitor captures and filters all these low-level operations before processing them into feature vectors, which are then forwarded to the File-based Behavioral Detector module of Minerva. To minimize overhead, we logically divide files into critical and non-critical: critical files encompass all user files that are high-priority targets for ransomware, while non-critical files comprise system and configuration files that can easily be recovered after an attack. The Disk Activity Monitor excludes any operation performed on non-critical files and computes per-file feature vectors for all other opened files.
4.1.2. File-Based Behavioral Detector.
Figure 1 illustrates the architecture of the File-based Behavioral Detector module of Minerva. This module comprises an ensemble of machine-learning classifiers trained to identify ransomware activity by analyzing the behavioral profile of files. For each critical file in the system, Minerva computes several features based on the operations received by the file within a predefined time window. These features define that file’s behavioral profile for that time window, which the Detector uses to determine whether the file has undergone malicious activity within the window. To withstand variations in ransomware behavior, such as adaptive attacks aimed at slowing ransomware activity to span multiple windows, Minerva employs a multi-tier architecture. Each tier implements a distinct fully-connected Deep Neural Network classifier trained on feature vectors computed over progressively longer time windows: low-tier classifiers utilize feature vectors computed over short time windows (e.g., $0.25$ sec), while high-tier classifiers utilize feature vectors calculated over longer windows (e.g., $4$ sec). At the end of each time window, Minerva queries the classifier of the corresponding tier to detect malicious behavioral profiles. In order to confirm malicious activity on a file, Minerva requires that at least one of its classifiers reports $K$ consecutive windows with detected malicious activity. If a classifier reports malicious activity for $K$ consecutive windows, all processes that were active on the specific file during those windows are flagged as ransomware processes. This procedure is carried out for each active critical file in the system that is receiving read or write operations.
The architecture of Minerva’s file-based detector is designed to accomplish two primary objectives. The first is to provide resilience to ransomware behavioral mutations aimed at slowing down encryption activity, which is achieved through Minerva’s tiered architecture. The second is to provide a configurable balance between detection time and false positive rates, which is accomplished by requiring $K$ consecutive windows with detected malicious activity. Section 6.2 elaborates on this trade-off.
4.2. Feature Analysis
Minerva leverages seven distinctive file-level features that accurately describe the behavioral profile of files. We carefully selected these features following an analysis of a dataset containing both benign and ransomware processes to validate our initial intuition behind the choice. Our experimental evaluation demonstrates the significant challenge ransomware faces in modifying its behavior to successfully evade all of these features, given the contrastive nature between different features.
4.2.1. Read - Write Data Mismatch.
The type of data read and written to a file is an important indicator of whether the file is seeing benign or malicious activity. For benign processes, the distribution of data read from a file and the distribution of data written to it are typically consistent. Conversely, ransomware tends to read non-encrypted data from files and write back encrypted data, exhibiting entirely distinct distributions. A sufficient approximation to capture this discrepancy can be obtained using entropy as a proxy for the distribution of data, as illustrated in Figures 2(a) and 2(e). These figures respectively plot the distribution and the CDF of the ratio between read and write entropy of operations performed by benign and ransomware. As we can see, benign file activity results in a Gaussian-like distribution centered around $1.0$, indicating that the read and write distributions roughly match. Ransomware, on the other hand, display a distribution that concentrates in the interval $[0,1]$, indicating higher entropy distribution for write operations compared to read operations. Similarly, the CDF highlights this same behavioral difference. While average operation entropy in isolation is not a robust metric (De Gaspari et al., 2022b), we demonstrate that, by incorporating it in our contrastive design, it provides sufficient and reliable information to aid in classification. We further discuss this in Section 4.3.
4.2.2. File Write Ratio.
Typically, benign processes write (or overwrite) small chunks of files over time, rarely rewriting an entire file. From the standpoint of file behavioral profiles, this implies that within a short time window, the number of distinct bytes written by all processes acting on the file should be significantly smaller than the file size. Conversely, to increase the probability of receiving a ransom, ransomware needs to completely overwrite as many files as possible, as swiftly as possible. Consequently, in a time window where ransomware activity is present, most or all of the file content will be overwritten. This behavioral asymmetry also holds for multiprocess ransomware that distributes write operations among multiple processes (De Gaspari et al., 2022a). Although each individual ransomware process writes only a fraction of the overall file, from the standpoint of file-level behavioral profiles, the combined activity of all ransomware processes results in a high write ratio. The percentage of a file written in a time window is, therefore, a defining, robust feature that helps characterize operations performed by both traditional and evasive multiprocess ransomware.
4.2.3. File Read Ratio.
While it is uncommon for a benign process to entirely overwrite a file, it frequently occurs that a benign process reads a file in its entirety. This behavior is prevalent for various user files, such as text documents, spreadsheets, and many others. Indeed, both ransomware and benign processes exhibit similar behaviors concerning the percentage of a file read, making it appear a weak feature for classification. However, its significance becomes evident when paired with the percentage of a file written. The primary techniques employed by evasive multiprocess ransomware to reduce feature expression are load splitting and function splitting: an individual ransomware process is divided into multiple processes, and both the number of operations and their type (i.e., read, write), are distributed among the processes. Splitting functionality among processes implies that, in certain cases, ransomware processes perform solely read operations on files, while sharing the read data with multiple writer processes through some stealth communication channel (De Gaspari et al., 2022a; Zhou et al., 2023). The writer processes then encrypt the data and overwrite the file at some later time. From the perspective of the file-based behavioral profile, this implies that there may be instances in which write operations of encrypted data occur within a time window, with no corresponding read operation.
Figures 2(d) and 2(h) highlight the fundamental difference between ransomware and benign processes behavior concerning file read/write ratio mismatch. We observe a substantial distinction between ransomware activity and benign process behavior both in the read/write ratio of files, thereby supporting our hypothesis that this feature offers valuable information for distinguishing between benign and ransomware activity in the system.
4.2.4. Number of Processes Reading or Writing the File.
Generally, in a time window with no ransomware activity, user files are accessed by a limited number of processes, often just one. This behavior also applies to certain traditional ransomware, where a single process reads and writes a file entirely. However, some traditional ransomware and all evasive multiprocess ransomware rely on the coordinated action of multiple processes targeting the same file, either to accelerate encryption or to evade detection. To imitate the behavior of benign processes, evasive multiprocess ransomware impose strict limits on the number of operations that each individual process is allowed to perform, as well as on the size of such operations (De Gaspari et al., 2022a; Zhou et al., 2023). This is achieved by spawning multiple distinct processes to read and write the file in rapid succession, thereby fully encrypting user files while satisfying these constraints. Consequently, the number of processes performing read or write operations on a file within a time window serves as a crucial feature for detecting both traditional and multiprocess ransomware.
Figures 2(c) and 2(g) illustrate the distribution and CDF of the number of processes acting on a file in a 1-second time window. We observe the inherent difference between ransomware and benign process behavior. The plots indicate that, in nearly all cases, only a single benign process accesses a file within a time window. While this behavior is also observed for some types of traditional ransomware, a clear distinction emerges for the remainder and for multiprocess ransomware.
4.2.5. Number of File Operations.
The number of read and write operations that a file undergoes within a given time window offers essential context to weigh the significance of the other features. Benign processes commonly execute occasional file operations of relatively small size over their lifetime, leading to file-based behavioral profiles characterized by a low number of operations within each time window. This is apparent from Figures 2(b) and 2(f), which respectively depict the distribution and CDF of the number of operations received by a file over 1-second intervals. Certain traditional ransomware favor using a limited number of large read and write operations during the encryption process, likely aiming to maximize I/O performance. Conversely, the majority of traditional ransomware utilize numerous small operations on the file, presumably to minimize the likelihood of detection. Similarly, in an effort to closely imitate the behavior of benign processes, evasive multiprocess ransomware imposes restrictions on the number of operations performed by each individual ransomware process and on the size of each operation (De Gaspari et al., 2022a; Zhou et al., 2023). The restriction imposed on individual operation size forces evasive multiprocess ransomware to use a greater overall number of read/write operations, distributing them across the various ransomware processes. In terms of file-based behavioral profiles, this behavior leads to an elevated number of file operations within a time window characterized by malicious activity, in contrast to one with only benign activity. These observations are validated by Figures 2(b) and 2(f), which highlight the propensity of ransomware to execute a substantially higher number of operations on a file within a given time window. Lastly, the number of operations on the file complements the number of processes feature, introducing another valuable dimension for the classifier to consider.
4.3. Adaptive Attack Resilience
When machine learning approaches are deployed in adversarial contexts, it is important to account for an adaptive adversary in the threat model, who alters their attack to evade the specific detector encountered. This section elaborates on the various types of adaptive attacks that may be crafted against Minerva, and outlines why our detector is resilient against them.
4.3.1. Evading the Read - Write Data Mismatch.
Minerva uses average operation entropy to identify mismatches between the distribution of data read from and written to a file. However, adversaries can easily manipulate the average entropy of operations, as demonstrated in literature (De Gaspari et al., 2020). Considering that the average entropy is computed across all operations within a time window, an adaptive ransomware attack simply requires injecting dummy low-entropy operations among the high-entropy encryption operations to evade this feature. These low-entropy operations can be executed by a single process for traditional ransomware or distributed across multiple processes for evasive multiprocess ransomware.
As we demonstrate in Section 6.5, Minerva is inherently resilient to this family of adaptive attacks. Indeed, arbitrary manipulation of the average entropy requires executing additional operations on the file within the same time window. While these extra operations achieve the intended goal of reducing average entropy, they also inadvertently increase the number of operations feature due to Minerva’s contrastive design, ultimately leading to detection.
4.3.2. Evading File Read and Write Ratios Features.
The file read and file write ratio features are directly correlated to the encryption speed of ransomware. The only method adaptive ransomware can adopt to alter these features is slowing down its encryption activity, leading to only a small portion of the file being read/written within a time window. However, this poses a significant challenge for ransomware, as its primary objective is to encrypt as many files as swiftly as possible before detection. If the ransomware were to slow down the encryption process excessively, it would render itself ineffective (Kharaz et al., 2016; Kharraz and Kirda, 2017).
We demonstrate in Section 6.5 that, due to its tiered architecture, Minerva effectively detects this family of adaptive ransomware, even when it employs substantial slowdowns of up to $80\%$.
4.3.3. Evading Number of Processes Reading or Writing the File Feature.
The number of processes feature is primarily addressed at detecting evasive multiprocess ransomware, which uses multiple independent processes to encrypt files. To evade this feature, adaptive ransomware would be required to reduce the number of processes operating on a file within a time window, thereby compromising its ability to distribute tasks across processes and imitate benign behavior. In turn, this leads the file-based profile of evasive multiprocess ransomware to resemble that of traditional ransomware. As shown in the experimental evaluation, Minerva excels at detecting this type of ransomware.
4.3.4. Evading Number of Operations Feature.
The number of operations feature primarily facilitates the detection of ransomware that attempts to minimize the size of its read/write operations. Both traditional and evasive multiprocess ransomware commonly adopt this strategy to make it more challenging for traditional detectors to identify their activity. Decreasing the number of operations within a time window requires a proportional increase in the size of individual reads/writes, thereby simplifying the detection task for traditional detectors such as antivirus.
Nonetheless, we demonstrate in Section 6.5 that Minerva is robust against this family of adaptive ransomware.
5. Experimental Setup
5.1. Dataset and Ransomware Defenses
5.1.1. Dataset of Process Operations.
The File-Based Behavioral Detector module of Minerva is trained on features calculated from multiple different datasets. The data for benign and traditional ransomware were obtained from the authors of ShieldFS (Continella et al., 2016) and RWGuard (Mehnaz et al., 2018), enriched with traces from other traditional ransomware families. Regarding the evasive multiprocess ransomware, we utilize the evasive multiprocess ransomware families presented by De Gaspari et al. (De Gaspari et al., 2020) in all their evasive configurations. Moreover, the dataset on adaptive ransomware was obtained by altering the behavior of the ransomware introduced by the same authors, to manipulate the features used by Minerva for detection, as discussed in Section 5.2.2.
5.1.2. Extracted Features Dataset
From the above-mentioned dataset, we derive four separate feature dataset: benign feature dataset, traditional ransomware feature dataset, evasive multiprocess ransomware, and adaptive ransomware feature dataset. Minerva models are only trained on benign, traditional, and multiprocess feature dataset. We stress that the adaptive ransomware features dataset is never used in the training process. It is only used in the testing phase, to assess the resilience of Minerva to adaptive attacks and its ability to generalize to unseen ransomware.
5.1.3. Ransomware Defenses
In our evaluation, we compare our approach against two state-of-the-art ransomware detection approaches: ShieldFS (Continella et al., 2016) and RWGuard (Mehnaz et al., 2018). As highlighted in recent publications (De Gaspari et al., 2022a; Zhou et al., 2023), ShieldFS and RWGuard are still the reference benchmarks for behavioral-based ransomware detection.
5.2. Implementation Details
5.2.1. Behavioral Detector
Minerva uses an ensemble multi-tier architecture, where each tier’s classifier monitors file activity occurring within a specific time window. In our implementation, we use the following time windows for each tier: $[0.25,0.5,1,2,4]$ seconds. These time windows were selected to balance detection speed, system efficiency, and robustness. For each tier, we employ a fully connected neural network classifier.
5.2.2. Adaptive Ransomware
To assess the resilience of Minerva against evasion attacks, we devise new families of adaptive ransomware specifically engineered to evade its detection. These adaptive ransomware are enhanced variants of the evasive multiprocess ransomware (De Gaspari et al., 2022a), improved to manipulate one or more of Minerva’s features to evade detection. We design three families of adaptive ransomware. (1) The adaptive R/W entropy ransomware adjusts the read and write entropy features by introducing artificial, low entropy read and write operations, thereby modulating the average entropy of operations within a given time window. We generate multiple variants of this adaptive family, with an up to $8$x reduction in average operation entropy. (2) The adaptive operation number ransomware alters the number of operations executed within a time window by increasing the size of each individual operation. We developed multiple variants of this family, resulting in up to $90\%$ reduction in number of operations. (3) The adaptive R/W ratio ransomware modifies the file read and write ratios within a specific window by slowing down the ransomware encryption activity. We create several variants for this family, with up to $80\%$ slowdown in the encryption speed.
We stress that all these adaptive families are not included in the training data, and are therefore unseen by the classifier.
Figure 3. True positive rate, true negative rate, and average detection time with standard deviation for varying K.
6. Evaluation
This section evaluates the performance of Minerva. We show that Minerva consistently detects different types of ransomware with low false positive rates, i.e., without impacting benign activity in the system. Moreover, we demonstrate that the fast detection time of our approach minimizes file loss rate, preventing widespread damage to the system. We evaluate Minerva on both unseen ransomware and unseen adaptive ransomware, which is crafted specifically to evade our detector. In both cases, we prove the generalization ability of Minerva and the robustness of our design. Lastly, we analyze the explainability of our classifier and provide insights on how Minerva distinguishes between benign and ransomware classes.
The evaluation is structured as follows. Section 6.1 introduces the evaluation metrics. Section 6.2 analyzes Minerva’s detection time. Sections 6.36.46.5 present the detection performance of Minerva over various types of ransomware. Section 6.6 provides insights on the explainability of Minerva and how different features affect the classification outcome. Lastly, Section 6.7 discusses the system overhead introduced by Minerva’s modules.
6.1. Metrics
As Minerva is a file-based detector, utilizing performance metrics based on the ratio of detected ransomware processes would be misleading. In each time window, Minerva constructs a behavioral profile for every opened user file in the system, and classification is conducted on a per-file basis within the window. Consequently, when evaluating Minerva’s performance, our focus lies on the ratio of detected files exhibiting malicious activity, rather than the ratio of detected ransomware instances. Evaluating Minerva based on ransomware detection rate would trivialize the detection task: as long as malicious ransomware activity on even a single file is detected, the detection performance would register as $100\%$. This metric not only lacks significance but also misrepresents the effectiveness of the defense. For instance, a ransomware that successfully encrypts most of the file system but is only detected during the encryption of the last few files would still be counted as a successful detection.
To better represent the effectiveness of our approach, we assess Minerva detection rate based on the number of files targeted by the ransomware that are successfully identified. We define True Positives (TP), True Negatives (TN), False Positives (FP), and False Negatives (FN) as follows.
  • TP: number of individual files exhibiting ransomware activity classified as malicious by Minerva.
  • TN: number of individual files without ransomware activity classified as benign by Minerva.
  • FP: number of individual files without ransomware activity classified as malicious by Minerva.
  • FN: number of individual files exhibiting ransomware activity classified as benign by Minerva.
Given the above definitions, we use the following standard performance metrics in our evaluation: True Positive Rate (TPR), True Negative Rate (TNR), accuracy, precision, recall, and F1 Score (F1).
When discussing our results, we denote the overall performance of the ensemble of Minerva models as stacked, while the performance of each individual Minerva classifier is referred to by the duration of its corresponding window (e.g., 0.25sec classifier is the classifier using 0.25sec of activity on a file to make a prediction).
Table 1. Performance analysis of Minerva against traditional and multiprocess ransomware attacks.
Benign Traditional Multiproc. Overall
Time window $TNR$ $TPR_{T}$ $TPR_{M}$ Accuracy Precision Recall F1
0.25sec 98.53 98.16 99.93 98.88 99.26 99.05 99.15
0.5sec 98.54 98.58 99.93 99.02 99.26 99.26 99.26
1sec 98.56 98.75 99.94 99.09 99.28 99.35 99.31
2sec 98.79 98.89 99.94 99.21 99.39 99.42 99.40
4sec 98.94 99.11 99.94 99.33 99.47 99.53 99.50
stacked 97.67 99.25 99.98 98.96 98.84 99.61 99.23
Figure 4. Cumulative Distribution Function of total ransomware activity duration on each file.
6.2. Detection Time
Detection time is a critical metric for assessing ransomware detection performance. From the ransomware perspective, hastening the malicious activity is paramount to achieving the objective of encrypting a maximum number of the victims’ files, thereby increasing the likelihood of ransom payment. Conversely, from the defender’s standpoint, minimizing the time required to detect the ongoing ransomware activity within the system reduces the probability of significant data loss. Furthermore, prompt detection of ransomware activity enables the adoption of near-zero overhead data recovery approaches, as we discuss in detail in Section 7.
In Figure 3, we report TPR, TNR, and the time to detect the ransomware activity in the system by Minerva for varying numbers of $K$. As mentioned in Section 4, $K$ indicates the number of consecutive windows where malicious activity is detected by Minerva in order to classify the operations as malicious. Different values of $K$ provide a trade-off to balance between detection time and false positive rates. Figure 3 shows that increasing the value of $K$ leads to a corresponding increase in $TNR$, indicating better end-user experience due to the decreased occurrence of benign processes flagged as ransomware. This is expected behavior, as for a well-trained classifier, the likelihood of $K$ consecutive misclassifications is inversely proportional to the value of $K$. Conversely, the average detection time is directly proportional to $K$. Figure 3 shows the proportionality relation is linear, with the average detection time being almost exactly equal to the minimum window size times $K$. This is due to the high TPR achieved by the 0.25sec window classifier, as we discuss later in Section 6.3. Finally, the overall TPR is nearly unaffected by $K$, decreasing by ${\sim}0.4$ percentage points between $K=1$ and $K=5$. The motivation behind this behavior is twofold. First, the TPR of Minerva is consistently high across all windows with malicious activity, meaning that typically every window with ransomware activity is correctly flagged by Minerva. Therefore, given $K$ windows with malicious activity, the probability that at least one is flagged as benign is low. Second, as illustrated in Figure 4, the total duration of ransomware activity on a single file lasts less than 1 second for ${\sim}95\%$ of all files. This behavior effectively puts an upper bound on the number of maximum windows $k$ required for detection. In fact, if there are less than $K$ windows of activity on a given file, Minerva will still attempt classification using the available windows. For example, considering the $0.25sec$ window classifier, since ${\sim}95\%$ of all files have less than 1 second of ransomware activity, Minerva will use at most $4$ windows ($1sec/0.25sec$) to perform classification on these files, regardless of $K$. Therefore, increasing $K$ beyond a certain value impacts the classification only on a very small percentage of files.
Based on this empirical evaluation, we found that $K=2$ provides the best trade-off between detection time, TNR, and TPR. For $K=2$, Minerva detects over 99% of the ransomware activity within $0.52sec$ ($0.25sec$ standard deviation) with low FPR, reducing the probability of significant data loss. Furthermore, as we discuss in Section 7, the fast detection rate of Minerva enables the use of near-zero overhead data recovery approaches, effectively nullifying data loss. Henceforth, unless explicitly stated otherwise, all subsequent experiments are presented for $K=2$.
6.3. Ransomware Detection Performance
This section evaluates the ransomware detection rate of Minerva and compares it to widely used benchmark papers in the area. We analyze the performance of our detector against two different categories of ransomware: traditional ransomware, and evasive multiprocess ransomware.
Table 2. Detection rate of ShieldFS, RWGuard, and Minerva against traditional and multiprocess ransomware. Results taken from (Continella et al., 2016; De Gaspari et al., 2022a).
Approach Traditional $TPR$ Multiprocess $TPR$
ShieldFS (Continella et al., 2016) 100.00 0.00
RWGuard (Mehnaz et al., 2018) 98.45 0.00
Minerva (ours) 99.25 99.98
Table 3. Performance analysis of Minerva against unseen ransomware families.
Benign Ransomware Overall
Time window $TNR$ $TPR$ Accuracy Precision Recall F1
0.25sec 98.49 97.73 98.11 98.48 97.73 98.11
0.5sec 98.45 98.44 98.45 98.45 98.44 98.45
1sec 98.41 98.62 98.52 98.41 98.62 98.52
2sec 98.65 98.79 98.72 98.66 98.79 98.72
4sec 99.09 98.92 99.00 99.08 98.92 99.00
stacked 97.59 99.07 98.33 97.63 99.07 98.35
6.3.1. Traditional Ransomware
Traditional ransomware comprises ransomware families that are used in real-world attacks and can be found in the wild. These families of ransomware typically exhibit low degrees of parallelism and spawn few processes to encrypt the target system and perform other malware-related activities, such as command-and-control communication (De Gaspari et al., 2020). Traditional ransomware sometimes employ multiple processes to speed up the encryption of the target system, but do not perform complex balancing of operations between the various processes to avoid detection (De Gaspari et al., 2022a).
Table 1 reports the performance of Minerva on the traditional ransomware test set. The true positive rate ($TPR_{T}$) value for all of the considered time windows is above $98\%$, reaching over $99\%$ on the $4sec$ time window. This means that all individual Minerva classifiers can accurately pinpoint traditional ransomware activity with less than 2% of false negatives. Similarly, the TNR ranges between ${\sim}98.5\%$ and ${\sim}99\%$ across all windows, highlighting the ability of Minerva to effectively detect ransomware without compromising user experience. Finally, the overall stacked performance of Minerva, which aggregates the predictions of each individual classifier to take a decision (see Section 4.1), reaches over $99\%$ TPR. The improved TPR of stacked Minerva compared to the $4sec$ window classifier derives from the independent nature of the errors made by each individual classifier, which results in higher overall detection rate.
The performance analysis in Table 1 emphasizes the ability of Minerva to generalize well across ransomware families, even when very short time windows are considered. This characteristic of Minerva is fundamental in ensuring rapid detection of ransomware activity ($0.52sec$ mean, see Section 6.2), which ensures low data loss rate and the compatibility of Minerva with near-zero overhead data protection techniques.
6.3.2. Evasive Multiprocess Ransomware
Evasive multiprocess ransomware families are characterized by their extensive use of parallelism, relying on multiple processes to encrypt the target system. These processes collaborate in a stealthy manner in order to perform complex balancing of the ransomware operations between multiple processes to avoid detection (De Gaspari et al., 2022a; Zhou et al., 2023). Some of these families leverage this covert inter-process collaboration to closely mimic benign process behavior (De Gaspari et al., 2020), entirely evading detection by traditional process-based behavioral detectors (Continella et al., 2016; Mehnaz et al., 2018; Kharraz and Kirda, 2017; Kharaz et al., 2016).
Table 1 reports the performance of Minerva on the evasive multiprocess ransomware test set. The test data is collected by running a variety of evasive techniques (De Gaspari et al., 2020), which are shown to completely evade behavioral-based ransomware detection (De Gaspari et al., 2022a). The true positive rate ($TPR_{M}$) is largely above $99\%$ for all of the time windows, which confirms our hypothesis that Minerva’s file-based behavioral profiling is resilient to this type of evasive multiprocess ransomware, contrary to prior works. We also note that there is essentially no performance variation between the different Minerva classifiers. While initially perplexing, this phenomenon can be explained upon closer analysis of the behavior of the multiprocess ransomware. By distributing the workload across multiple processes, the activity of the ransomware on each individual file typically lasts less than $0.5sec$ — the minimum time required for detection by Minerva with $K=2$. Consequently, all Minerva classifiers see the full ransomware activity on the file, which explains the negligible performance variations between different windows. Furthermore, as elaborated later in Section 6.6, the variability in behavior among the various processes of multiprocess ransomware is minimal compared to traditional ransomware, which aids in clarifying the near-perfect detection performance.
6.3.3. Comparison with SOTA
We further compare the performance of Minerva against well-known benchmark ransomware detection approaches (Continella et al., 2016; Mehnaz et al., 2018). Table 2 presents our comparison. As illustrated in the table, Minerva achieves performance comparable to state-of-the-art approaches against traditional ransomware. Furthermore, Minerva successfully detects evasive multiprocess ransomware, a feat that existing approaches fail to accomplish.
(a) Detection rate of adaptive ransomware relying on reducing the number of operations for different reduction factors, K=2.
(b) Detection rate of adaptive ransomware relying on reducing the read/write frequency for different slowdown factors, K=1.
(c) Detection rate of adaptive ransomware relying on reducing the read/write frequency for different slowdown factors, K=2.
Figure 5. Adaptive ransomware detection rate, for different adaptive approaches and different configurations.
6.4. Unseen Detection Performance
Our prior results demonstrate that Minerva exhibits excellent performance in detecting known ransomware. Nevertheless, given the continual emergence of new ransomware variants and the ongoing evolution of existing ones, an arguably more critical evaluation criterion for ransomware detectors lies in their capability to generalize to unseen data.
In line with previous works (Continella et al., 2016; Mehnaz et al., 2018; Kharraz and Kirda, 2017), we assess Minerva’s capability to generalize to unseen ransomware by removing $10\%$ of the variants from our training data and evaluating detection performance on these withheld samples. This evaluation provides a measure of the expected predictive and generalization capabilities of Minerva on future, unknown ransomware variants. Table 3 reports the results of our test. As highlighted in the table, Minerva’s performance on unseen variants closely aligns with its performance on known variants, exhibiting only a minor performance degradation of ${\sim}0.2$ percentage points across all classifiers. These findings underscore the ability of Minerva to learn descriptive features that accurately capture the diverse range of behaviors displayed by ransomware.
Table 4. Average stacked detection rate of Minerva for different adaptive ransomware families.
Adaptive Type $TPR$ Mean Det. Time (sec) Det. Time Std. Dev.
R/W Entropy 99.93 0.50 0.02
R/W Ratio 99.94 1.82 1.07
Op. Num. 98.16 0.56 0.45
6.5. Adaptive Detection Performance
As previously discussed, our primary objective with the design of Minerva is to develop a robust classifier, capable of withstanding adaptive attacks specifically engineered to evade its detection. As outlined in Section 4, we developed Minerva following a contrastive design, where altering one feature triggers changes to one or more other features. This characteristic renders evasion attempts by adaptive ransomware precarious, as any alteration intended to evade detection inadvertently induces changes in other features used by Minerva, ultimately leading to detection.
We assess the effectiveness of our contrastive design by evaluating Minerva against ransomware that has been specifically altered to evade Minerva detection mechanisms. This adaptive ransomware is an enhanced variant of the evasive multiprocess ransomware (De Gaspari et al., 2022a), improved to manipulate one or more of the features used by Minerva to circumvent detection, as elaborated in Section 5.2.2. We stress that all these adaptive families are not included in the training data, and are therefore unseen by the classifier. Table 4 illustrates the stacked detection rate of Minerva for different families of adaptive ransomware. Notably, the TPR remains largely unaffected by the behavioral modifications of the different adaptive families. Specifically, adaptive ransomware targeting the read-write entropy features are promptly detected by Minerva within $0.5sec$, which is the minimum detection time when $K=2$. Similarly, adaptive ransomware altering read- and write-ratio behaviors is also consistently detected. However, the mean detection time for this adaptive family is significantly extended (${\sim}3$x), with a $1sec$ standard deviation. These findings are expected, as this adaptive ransomware family reduces the R/W ratio feature by aggressively slowing down encryption activity (up to $80\%$ slowdown, as outlined in Section 5.2.2). Nevertheless, owing to the contrastive design of Minerva, detection performance remains excellent, as the classifiers leveraging long-term windows effectively identify the slow encryption activity. The sole adaptive family achieving limited success in evading Minerva are the adaptive operation number ransomware. These ransomware reduce the number of read and write operations by increasing the size of each individual operation. However, as highlighted in Table 4, Minerva continues to deliver excellent detection performance against this adaptive family. Furthermore, since this evasion strategy entails reducing the number of operations while increasing their size, attempting to evade Minerva using this method renders the attack easily detectable through process-based detection (De Gaspari et al., 2022a). Finally, we also evaluated Minerva against compatible combinations of evasive strategies (r/w entropy + r/w ratio, r/w ratio + op. num.), obtaining similar detection results.
(a) SHAP values on 1 second window, benign process features.
(b) SHAP values on 1 second window, traditional ransomware.
(c) SHAP values on 1 second window, multiprocess ransomware.
(d) SHAP values on 1 second window, adaptive entropy ransomware.
(e) SHAP values on 1 second window, R/W ratio adaptive ransomware.
(f) SHAP values on 1 second window, Operation Number adaptive ransomware.
Figure 6. SHAP-based analysis of the Minerva features on the detection of benign and different ransomware types considered (i.e., traditional, multiprocess and adaptive).
In Figure 5, we further study the detection rate distribution across Minerva’s classifiers for the different adaptive families. We exclude from this analysis r/w entropy adaptive family, as it is always detected by the short-term $0.25sec$ classifier. Figure 5(a) plots the detection rate against adaptive ransomware that manipulates the number of operations feature. From the figure, we notice that the vast majority of all the family’s variants ($>90\%$) are detected within $0.5sec$ by the short-term $0.25sec$ window when $K=2$. These results are expected, as this adaptive family does not attempt to slow down encryption activity, and therefore we expect a detection time in line with the overall mean detection time (see Figure 3). Figures 5(b) and 5(c) depict the detection rate distribution for the adaptive R/W ratio family of ransomware for $K=1$ and $K=2$ respectively. As expected, the detection distribution among the Minerva classifiers significantly differs compared to other ransomware families, with a tendency towards longer-term classifiers as the slowdown ratio increases. Furthermore, interesting insights can be derived by comparing the detection distribution for $K=1$ and $K=2$. Initially, one might expect the detection distribution for $K=2$ to closely resemble that of $K=1$, albeit with double the detection time. However, we observe that this expectation is incorrect. In reality, the minimum detection time for adaptive R/W ransomware increases fourfold between the two settings, while the maximum detection time increases twofold, as initially anticipated. This behavior is explained by the weaker detection performance of Minerva’s short-term classifiers on individual windows of ransomware activity for this adaptive family. While the short-term term classifiers effectively detect at least one window of malicious activity (sufficient for detection when $K=1$), they fail to detect two consecutive windows when the ransomware activity slows down. This ultimately leads to a bias in the detection distribution towards long-term classifiers, and, therefore, longer detection times.
6.6. Feature contribution to decision-making
Explainability is a complex yet crucial aspect in the deployment of DNNs, especially in critical domains such as cybersecurity. While previous sections have empirically showcased the effectiveness of Minerva, its ability to generalize to unseen data, and its resilience to adaptive evasion attacks, it is also important to provide insights into the reasons behind Minerva’s effectiveness. In this section, we aim to bridge this gap by examining the contributions of each of Minerva’s features to the final classification outcome. We leverage the SHapley Additive exPlanations (SHAP) (Lundberg and Lee, 2017) technique, a widely adopted method to interpret model predictions. SHAP values are based on a game theoretic approach to assign importance values to each feature in a model. Features with positive SHAP values influence the prediction towards the positive class, whereas those with negative SHAP values steer the prediction towards the negative class. The magnitude of each individual feature indicates the strength of its contribution to the final prediction. As a reminder, Minerva is a file-based behavioral classifier. Therefore, it is understood that the ensuing discussion on feature importance and behavioral profiles is conducted from the perspective of activity on individual files.
Figure 6 presents the SHAP plots for the $1sec$ window Minerva classifier, highlighting the importance of each selected feature to the final classification and how feature importance changes for different types of ransomware. Each plot presents the SHAP values of Minerva’s features for 100 randomly sampled data points. For ransomware, we generate separate SHAP plots for each type, allowing us to highlight the distinctive contribution of features to classification, based on the diverse behavior exhibited by each ransomware type. Figures 6(a)6(b), and 6(c) display the individual contributions of the Minerva features to the prediction of benign, traditional ransomware and multiprocess ransomware classes for the one-second window model. An immediate observation is the significance of the write ratio feature as a distinguisher between benign and ransomware instances. This finding is expected, as benign applications typically perform partial writes to files, while ransomware aims to fully rewrite user files as part of its malicious activity. A similar pattern emerges with the read ratio feature, where benign applications often read only portions of files, while ransomware tend to read files fully to encrypt them. Figure 6(b), however, reveals exceptions to this trend, indicating that some traditional ransomware variants also exhibit partial file read and write behaviors. We hypothesize that these ransomware variants prioritize speed over completeness of encryption, sacrificing full file encryption to achieve higher overall encryption speed. Other important features are operation number and process number. Figure 6 underscores the descriptive nature of these features, as ransomware typically exhibits a notably higher number of operations and number of processes acting on files during the encryption activity, as previously highlighted in Figures 2(b) and 2(c). This characteristic is especially prominent for multiprocess ransomware, where multiple processes collaborate in the encryption process to mimic the per-process behavior of benign processes. Finally, the read and write entropy features play a crucial role in discriminating between benign and ransomware activity, with the write entropy feature having a particularly strong influence on the model’s classification for ransomware. Conversely, the impact of these features on benign activity is ambiguous. This outcome is expected, as certain benign applications display read/write entropy profiles akin to ransomware, as previously outlined in Section 4.2.
Figures 6(e),and 6(f) illustrate the SHAP plots for adaptive R/W ratio and adaptive operation number, respectively. We observe that these plots exhibit minimal variation when compared to the non-adaptive multiprocess ransomware depicted in Figure 6(c). These findings are expected, as the behavioral changes introduced by these adaptive ransomware families have limited impact on individual features of Minerva. Specifically, for adaptive R/W ransomware, the behavioral changes typically affect only short-term classifiers, whose detection windows are too brief to capture a meaningful portion of the malicious behavior, resulting in classification failure. Conversely, for long-term classifiers, the entirety (or a significant portion) of the malicious activity is captured, leading to a behavioral profile — and hence a SHAP profile — closely resembling that of non-adaptive families. Likewise, adaptive operation number ransomware primarily alter the number of operations executed during encryption activities. This behavior is evident in Figure 6(f), where the contribution of the operation number feature on the final classification is negligible, whereas the contribution of the other features remains largely unchanged. Lastly, Figure 6(d) clearly illustrates the influence of contrastive features such as entropy and number of operations on classification outcome. We observe that while the entropy features exhibit a minimal to negative impact on ransomware classification, the contribution of the operation number features shows a considerable increase. This behavior is expected, as adaptive entropy ransomware relies on dummy, low entropy operations to decrease the average read/write entropy and evade detection.
At a higher level, Figures 6(a)6(b), and 6(c) provide insights on the detection performance of Minerva discussed in Section 6.3. Specifically, we observe that the SHAP plot for benign processes — and thus their behavioral profiles — display a greater degree of variability compared to ransomware, making it challenging for Minerva to learn a general representation for this class. Conversely, ransomware, particularly multiprocess ransomware, exhibit more uniform behavioral profiles, thereby simplifying the task of learning a suitable representation for classification.
6.7. Minerva’s detection module overhead
6.7.1. Disk activity monitor overhead
The Disk Activity Monitor module is a shim providing system-wide, real-time information on all file system operations performed on critical files (see Section 4.1.1). Minerva, similar to previous research in this domain (Continella et al., 2016; Mehnaz et al., 2018) relies on an I/O Request Packets Logger driver (IRPLogger) to capture file system activity of interest. This collected information is further processed to generate feature vectors utilized by Minerva for detection. Consequently, the overhead induced by the disk activity monitor primarily stems from the continuous operation of the IRPLogger. Previous in-depth analyses indicate that, on average, the IRPLogger incurs a $1.02\%$ increase in the system’s CPU usage (Mehnaz et al., 2018), with no notable performance degradation from the users’ standpoint (Continella et al., 2016). Additionally, it is important to highlight that Minerva maintains behavioral profiles exclusively for opened user files, further minimizing its impact on the system.
6.7.2. Sample prediction time
Minerva relies on fully-connected DNN classifiers to determine whether a file activity is benign or malicious. To provide an overview of the time overhead associated with Minerva’s prediction process, we invoked the predict functionality of Minerva’s detector on $1,000$ individual samples and recorded the timing. On average, the Minerva’s classifiers take $0.0138sec$ to predict a single sample ($0.0016sec$ standard deviation), introducing negligible impact on the system and Minerva’s decision-making pipeline.
7. Discussion and Limitations
As outlined in Section 6, Minerva shows significant effectiveness in promptly identifying ransomware activity within the system. Depending on the encryption method adopted by the ransomware, however, there exists a potential for data loss. Our evaluation indicates that, for $K=2$, Minerva achieves an average detection time of $0.52sec$. This corresponds to approximately 19 files encrypted before detection, on average (Splunk, 2022). Nonetheless, certain ransomware families such as LockBit and Babuk demonstrate much higher speeds, capable of encrypting up to ${\sim}130$ files within the same time frame (Splunk, 2022). Furthermore, in theory, ransomware could generate a separate process for each file in the system to execute the encryption process, which would result in complete data loss. While such an approach would be impractical and quickly detectable through alternative means, sophisticated ransomware could employ a simplified version of this encryption method to maximize data loss. We emphasize that this concern is not exclusive to Minerva, but is a shared consideration for all ransomware detection systems.
Owing to its rapid detection rate, Minerva enables the utilization of near-zero overhead data recovery methods to effectively eliminate the risk of data loss. Modern operating systems employ a technique known as I/O buffering, which utilizes a page cache to temporarily store data in main memory before writing it to an I/O device. Operating systems periodically flush portions of the page cache: Windows flushes $1/8th$ of the cache every second (Russinovich et al., 2012), while Linux, by default, holds pages for a maximum of 30 seconds. Given that Minerva’s detection rate largely falls below the cache flushing rate of both operating systems, it becomes feasible to seamlessly adopt cache-based recovery techniques such as (Elkhail et al., 2023). These techniques modify cache flushing policies to prevent content flagged as malicious in the cache pages from being written to disk, providing a near-zero overhead data recovery process.
Lastly, another limitation of Minerva lies in its coarse-grained detection. Minerva detects ransomware activity on a per-file, per-window basis. While this allows Minerva to detect complex attacks such as evasive multiprocess ransomware, it also limits the detection granularity at the file, rather than at the process level. Thus, when Minerva detects ransomware activity on a file during a given time window, all processes that were acting on the file during the window are classified as ransomware and suspended. This group may include benign processes, that wrote data on the file within the same window as the ransomware. We argue, however, that this is a minor limitation. First, operating systems prevent concurrent writes on the same file. Therefore, while the ransomware is encrypting a file, no benign write activity can happen. Second, since Minerva detection time is very low, even if benign activity happened in the same window as ransomware, on average, only $0.5sec$ of benign activity would be lost. Lastly, we argue that the benefits of promptly stopping ransomware activity, outweigh the disadvantage of dropping a very short amount of benign user activity.
8. Conclusions
We presented Minerva, a novel approach to ransomware detection and mitigation. In contrast to existing detection methods, Minerva constructs behavioral profiles of files based on all the operations they receive within a specific time window. Thanks to its contrastive design approach, Minerva’s detection remains robust against complex multiprocess ransomware and adaptive ransomware specifically engineered to evade detection. Through a comprehensive evaluation, we demonstrated that Minerva is highly effective against all types of known ransomware attacks, unseen ransomware attacks, as well as evasive variants designed to bypass Minerva, underscoring the efficacy of our approach. Lastly, Minerva exhibits rapid ransomware activity detection, enabling the adoption of near-zero overhead data loss prevention techniques. We believe that these findings represent a significant advancement in the development of robust and resilient countermeasures against the escalating ransomware threat landscape.
9. Acknowledgments
This work was supported by the PON program of the Italian MUR under the project “Application of Machine Learning to improve olive yield and reduce climate change impact”, project SERICS (PE00000014) under the NRRP MUR program funded by the EU-NextGenerationEU, and project AutoAD: Using Active Defense to Defeat Cyber Adversaries funded by Sapienza University of Rome.